# Driving as a Diagnostic Tool: Scenario-based Cognitive Assessment in Older Drivers From Driving Video

### v1

@@ 2.2.1 Scenario-specific Video Sample
The overall pipeline for scenario-specific video sample preparation is illustrated in Fig. 4. Let $\mathcal{D}={D_{1},D_{2},\dots,D_{d}}$ be the set of subjects that participated in this study. Each subject $D_{i}$ performs a set of drives {${dr_{1},dr_{2},\dots,dr_{n}}$} where $n$ is the maximum number of drives per subject. Each drive is further segmented into short video clips ${V_{1},V_{2},\dots,V_{m}}$, where $m$ is the maximum number of video clips per drive. These video clips are recorded under different driving scenarios $\mathcal{S}={Sc_{1},Sc_{2},\dots,Sc_{s}}$. From these clips, a subset of scenario-specific video samples $V_{k}^{(D_{i},Sc_{j})}$ is filtered for each subject-scenario pair $(D_{i},Sc_{j})$ where $k$ is the index of the sampled video clip. This resulted in a quality dataset of labeled video samples that were used as inputs to the large vision model in the next stage. Additionally, the hierarchical rollout of video samples across subjects, drives, and scenarios ensures a structured and scalable representation across varying cognitive profiles and driving contexts.
2.2.2 Large Vision Model-based Framework
We feed the video samples into a pre-trained Large Vision Transformer model [17] to perform the video-to-vector generation. The extracted vector representations are stored in Amazon AWS Cloud for further analysis. The goal of this framework is to transform raw driving videos into discriminative vector representations for downstream classification of cognitive aging groups. This is achieved via the following four-step pipeline: The framework follows four main steps: The entire framework includes four steps as follows:
Step 1: Video-to-embedding Generation
@@ 2.2.2 Large Vision Model-based Framework
Each video sample $V_{k}^{(D_{i},Sc_{j})}$ is processed through a pre-trained large vision model $F_{\theta}$ to obtain a fixed-size vector space:
@@
where $Z_{k}$ represents the vector representation of each video sample $V_{k}^{(D_{i},Sc_{j})}$ in an $m$-dimensional latent space.
@@
To interpret the structure of the vector space and the separability of the aging groups in each scenario setting, we apply a dimensionality reduction function $g:\mathbb{R}^{m}\rightarrow\mathbb{R}^{n}$ (e.g., PCA [18], t-SNE [19]) to obtain:
$\tilde{Z}_{k}=g(Z_{k}),\quad\tilde{Z}_{i}\in\mathbb{R}^{n},\text{ where }n\in\{50,100,200,\dots\}$
This step is performed in a way to preserve the majority of the information embedded in the high-dimensional video representation while mapping it into a lower-dimensional vector space. It allows us to observe how the visual features related to the driving behavior in different scenarios are clustered in a low-dimensional space and select the prominent scenarios for the next steps. As shown in Fig. 5, a demo reduced 2D $(n=2)$ representations are plotted to visualize group separability within each scenario.
@@
To evaluate the relevance of a scenario $Sc_{j}\in\mathcal{S}$ for classification, we compute a pairwise distance metric (e.g., centroid, Euclidean) over the embeddings belonging to different classes (e.g., Normal, AD). The scenario with the highest inter-group distance is selected:
@@
$Sc^{*}=\arg\max_{Sc_{j}\in\mathcal{S}}\ \text{Dist}_{\text{avg}}(Z_{\text{class}_{normal}}^{(Sc_{j})},Z_{\text{class}_{AD}}^{(Sc_{j})})$
where $Z_{\text{class}_{normal}}^{(Sc_{j})}$ and $Z_{\text{class}_{AD}}^{(Sc_{j})}$ denote the set of embeddings under scenario $Sc_{j}$ for class Normal and AD respectively.
Step 4: Classification using RF Model
The scenario-specific embeddings are then used to train a Random Forest (RF) [20] classifier $H:\mathbb{R}^{m}\rightarrow\mathcal{Y}$, where $\mathcal{Y}$ is the set of class labels:
$\hat{Y}_{k}=H(Z_{k})$
Final model performance is assessed using metrics such as $\mathbf{(a)}$, precision$\mathbf{(P)}$, recall$\mathbf{(R)}$ and $\mathbf{(F1)}$ under multiple sampling strategies (e.g., random vs. driver-level separation) to ensure generalizability. We report average classification accuracy under two sampling strategies:
@@
    Driver-level Separation (DLS): To evaluate the generalizability of the proposed framework across different individuals, we employed a leave-$k$-drivers-out cross-validation strategy [21], where $k$ drivers are excluded from the training set and reserved exclusively for testing. This procedure is repeated over $r$ independent runs, each with a different subset of left-out drivers. Finally, the results across these runs are averaged to get the performance metrics $\mathbf{(a,P,R,F1)}$. It ensures robustness and mitigates variance due to subject-specific biases.
This setup allows us to evaluate both intra-subject consistency and inter-subject generalization in the learned representations.
Algorithm 1 Scenario-Based Video Analysis
1:Input: Set of video samples $\mathcal{V}={V_{k}^{(D_{i},Sc_{j})}}$ for subject $D_{i}$ in scenario $Sc_{j}$
2:Output: Trained classifier $H$, scenario relevance ranking ${\text{Dist}_{\text{avg}}(Sc_{j})}$
3:for each scenario $Sc_{j}\in\mathcal{S}$ do
4:     Initialize $\mathcal{E}_{Sc_{j}}\leftarrow\emptyset$triangleright Store embeddings
5:     for each subject $D_{i}\in\mathcal{D}$ do
6:         for each video $V_{k}^{(D_{i},Sc_{j})}$ do
7:              $Z_{k}\leftarrow F_{\theta}(V_{k}^{(D_{i},Sc_{j})})$triangleright Video-to-vector
8:              $\mathcal{E}_{Sc_{j}}\leftarrow\mathcal{E}_{Sc_{j}}\cup{Z_{k}}$
9:         end for
10:     end for
11:     ~$\tilde{\mathcal{E}}_{Sc_{j}}\leftarrow g(\mathcal{E}_{Sc_{j}})$triangleright Dimensionality reduction, $g:\mathbb{R}^{m}\to\mathbb{R}^{n}$
12:     Compute $\text{Dist}_{\text{avg}}(Sc_{j})\leftarrow\text{Distance}(\tilde{\mathcal{E}}^{\text{normal}},\tilde{\mathcal{E}}^{\text{mci}})$
13:end for
14:$Sc^{*}\leftarrow\arg\max_{Sc_{j}}\ \text{Dist}_{\text{avg}}(Sc_{j})$triangleright Most discriminative scenario
15:Train RF classifier $H:\mathbb{R}^{m}\rightarrow\mathcal{Y}$ using $\mathcal{E}_{Sc^{*}}$ and corresponding labels
16:return $H$, scenario ranking ${\text{Dist}_{\text{avg}}(Sc_{j})}$
@@ 3.1 Scenario-based Roadway Selection
To investigate the relationship between driving behavior and cognitive status, we selected two distinct types of roadway scenarios where older drivers commonly exhibit prominent visual or maneuver-based patterns. These patterns can serve as interpretable behavioral cues for identifying cognitive decline. The selected driving scenarios for our video-based analysis are:
@@
We utilize Omaha most transited route database for filtering the roadway scenarios. From different federal functional categories (e.g., Local, Major collector, Minor collector etc.), we observe and select the most prominent scenario combination for this study. Figure 6 shows the GIS map highlighting the two roadway scenarios for our video analysis. We considered a single driving video having $\leq 0.31$ mile or $\leq 500$ meter length as a route segment. Omaha most transited route database consists of 24,106 route segments. In total, we filter out $964$ unique route segments, consisting of $632$ freeway-interchange segments (highlighted in green in Figure 6(b)) and $332$ interstate segments (highlighted in blue in Figure 6(a)). Table 2 summarizes the key statistics – a balance of the selected data and driving exposure of the cognitive groups. Additionally, we selected the route segments that maximize the number of drives for both cognitive groups. The avg. trips per unique subject column indicates that each driver has driven on the chosen roadway types at least 3.5 times, ensuring sufficient behavioral representation per subject.
To interpret the driving exposure data distribution, we define several key variables below:
@@
    $M_{u}$: Number of unique trips by the subjects with MCI
@@
    $A_{u}$: Number of unique trips by the subjects with AD
@@
    $T_{u}=N_{u}+M_{u}+A_{u}$
    It does not include repeated trips which means multiple driving sessions recorded on the same drive segment by the same individuals
@@
    $T_{avg}$: Average number of trips per subject in a specific drive segment. It gives an approximation that drivers, on average, drove more on Interstate than on Freeway interchanges.
Table 2: Trip details by roadway types
Scenario Normal MCI AD Total Total Avg. trips per
categories trips($N_{u}$) trips($M_{u}$) trips($A_{u}$) trips($T_{u}$) trips($\overline{T}_{u}$) subject($T_{avg}$)
@@
$N_{u}$: unique trips by Normal group; $M_{u}$: unique trips by MCI group; $A_{u}$: unique trips by AD group; $T_{u}$: total unique trips; $\overline{T}_{u}$: total non-unique trips; Fwy-int.: Freeway interchange.
We leveraged “unique drives” vs. “repeated drives” per subject across all the route segments to indicate statistical reliability and diversity in driving exposure by the two cognitive groups. The average trips per subject suggest that each participant had substantial driving trips with the selected roadway scenario with a significant amount of video segments for our analysis.
Figure 7 shows a boxplot distribution of speed limits and the route distance for both scenarios. As we observe, the freeway-interchange segments tend to have lower speed limits and shorter distances, which works as a transitional driving scenario, involving challenging driving behavior and fast decision-making. In contrast, the interstate segments exhibit higher speed limits and longer travel distances, which indicate an uninterrupted highway driving scenario with relaxed driving behavior.
Figure 7: Boxplot distributions of the speed limit and route distance for Interstate and Freeway-interchange roadways. The middle line shows the median speed limit and median route distances.
@@
    Fwy-interchange scenario, $S_{\text{fwy}}=37.5$ mph;
@@
    Fwy-interchange scenario, $D_{\text{fwy}}=0.12$ miles;

### v2

@@ 2.3 Scenario-specific Video Sampling
The overall pipeline for scenario-specific video sampling is illustrated in Figure 4. Let $\mathcal{D}={D_{1},D_{2},\dots,D_{d}}$ be the set of subjects that were filtered out in section 2.2. Each subject $D_{i}$ performed a set of drives $dr_{1},dr_{2},\dots,dr_{n}$}} where $n$ is the maximum number of drives per subject. Each drive was further segmented into short video clips $V_{1},V_{2},\dots,V_{m}$}, where $m$ is the maximum number of video clips per drive. These video clips were recorded under different driving scenarios, $\mathcal{S}={Sc_{1},Sc_{2},\dots,Sc_{s}}$. From these video clips, a subset of scenario-specific video samples $V_{k}^{(D_{i},Sc_{j})}$ was filtered for each subject-scenario pair $(D_{i},Sc_{j})$ where $k$ is the index of the sampled video clip, $i$ is the index of the subject and $j$ is the index of the scenario. This resulted in a quality dataset of labeled video samples that were used as inputs to the large vision model in the next stage. This sampling strategy ensures sufficient data representation across both aging groups (Normal vs. MD) and allows the proposed method to be validated on balanced and identical driving scenarios. Additionally, the hierarchical video sampling steps across subjects, drives, and scenarios ensure a structured and scalable representation across varying cognitive profiles and driving contexts.
2.4 Large Vision Model-based Framework
The video samples derived in section 2.3 were fed into a pre-trained large Vision Transformer model [37] to generate vectorized representations of the video data. The objective of this framework is to encode driving videos into discriminative vector representations for downstream classification of cognitive aging groups. These representations capture prominent visual features that contrast the driving behavior of the two cognitive groups and distinguish them based on the vector’s distance in feature space (illustrated in Figure 5). The entire framework includes four steps as follows:
Step 1: Video-to-vector generation
@@ 2.4 Large Vision Model-based Framework
Each video sample $V_{k}^{(D_{i},Sc_{j})}$ is passed through a pre-trained large vision model $F_{\theta}$ to obtain a fixed-size vector space:
@@
where $Z_{k}$ denotes the vector representation of each video sample $V_{k}^{(D_{i},Sc_{j})}$ in an $m$-dimensional latent space.
@@
To interpret the structure of the vector space and the separability of the aging groups in each scenario setting, we apply PCA [38], a dimensionality reduction function $g:\mathbb{R}^{m}\rightarrow\mathbb{R}^{n}$ to obtain:
$\tilde{Z}_{k}=g(Z_{k}),\quad\tilde{Z}_{i}\in\mathbb{R}^{n},\text{ where }n\in\{50,100,200\},n\ll m$
This step was performed in a way to preserve the majority of the information embedded in the high-dimensional video representation while mapping it into a low-dimensional vector space. It allows us to observe how the visual features related to the driving behavior in different scenarios for the two aging groups are clustered in a low-dimensional feature space.
@@
To evaluate the relevance of a scenario $Sc_{j}\in\mathcal{S}$ for distinguishing the two aging groups, we computed a class-wise Euclidean distance in the reduced feature space. In Figure 5, an example of a dimension-reduced representation ($n=2$) with the group separability in feature space is illustrated. The Euclidean distance between the class centroids was obtained as follows:
$\boldsymbol{\mu}_{\text{normal}}^{(Sc_{j})}=\frac{1}{|\tilde{Z}_{\text{normal}}^{(Sc_{j})}|}\sum_{\mathbf{z}\in\tilde{Z}_{\text{normal}}^{(Sc_{j})}}\mathbf{z},\quad\boldsymbol{\mu}_{\text{MD}}^{(Sc_{j})}=\frac{1}{|\tilde{Z}_{\text{MD}}^{(Sc_{j})}|}\sum_{\mathbf{z}\in\tilde{Z}_{\text{MD}}^{(Sc_{j})}}\mathbf{z}$
Here, $\tilde{Z}_{\text{normal}}^{(Sc_{j})}$ and $\tilde{Z}_{\text{MD}}^{(Sc_{j})}$ denote the sets of low-dimensional vectors under scenario $Sc_{j}$ for the Normal and MD-aging groups respectively. Each centroid $\boldsymbol{\mu}_{\text{normal}}^{(Sc_{j})},\boldsymbol{\mu}_{\text{MD}}^{(Sc_{j})}\in\mathbb{R}^{n}$ represents the average position of the corresponding group in the feature space for the same scenario. The Euclidean distance between the two centroids was then obtained by:
$\mathrm{Dist}(Sc_{j})=\left\|\boldsymbol{\mu}_{\text{normal}}^{(Sc_{j})}-\boldsymbol{\mu}_{\text{MD}}^{(Sc_{j})}\right\|_{2}$
Then, $\mathrm{Dist}(Sc_{j})$ is used to determine the most discriminative scenario as:
$Sc_{\max}=\arg\max_{Sc_{j}\in\mathcal{S}}\;\mathrm{Dist}(Sc_{j})$
where $Sc_{\max}$ denotes the most discriminative scenario that maximizes the inter-group distance between the Normal and MD-aging groups centroids in the feature space.
Step 4: Validation of Scenario Selection
To validate the scenario selection, two identical Random Forest (RF) [39] classifiers were trained and evaluated on the most and least discriminative scenarios obtained in Step 3. In addition to the most discriminative scenario, we calculated the least discriminative scenario as:
$Sc_{\min}=\arg\min_{Sc_{j}\in\mathcal{S}}\;\mathrm{Dist}(Sc_{j})$
Let $\tilde{\mathcal{Z}}^{(Sc)}=\{\tilde{Z}_{k}^{(Sc)}\}$ denotes the reduced dimentional vectors (from Step 2) and $\mathcal{Y}^{(Sc)}=\{Y_{k}^{(Sc)}\}$ their corresponding labels for scenario $Sc\in\{Sc_{\max},Sc_{\min}\}$. We trained two RF classifiers with identical hyperparameters on the reduced feature space for both scenarios $\{Sc_{\max},Sc_{\min}\}$:
$\hat{Y}_{k}^{(Sc)}\;=\;\tilde{H}^{(Sc)}\!\left(\tilde{Z}_{k}^{(Sc)}\right),\qquad\tilde{H}^{(Sc)}:\mathbb{R}^{n}\rightarrow\mathcal{Y},$
[…]
$\bigl{\{}\mathbf{a},\,\mathbf{P},\,\mathbf{R},\,\mathbf{F1}\bigr{\}}_{Sc_{\max}}\quad\text{vs.}\quad\bigl{\{}\mathbf{a},\,\mathbf{P},\,\mathbf{R},\,\mathbf{F1}\bigr{\}}_{Sc_{\min}}$
and reported the performance difference as a validation criterion. A consistent performance gain of $\tilde{H}^{(Sc_{\max})}$ over $\tilde{H}^{(Sc_{\min})}$ validates the proposed scenario-based approach by demonstrating that scenarios with larger inter-centroid separation yield stronger discriminability between the two cognitive aging groups. The entire framework is shown in Algorithm 1. Moreover, we further evaluated the validity of the framework under different data sampling strategies. Multiple sampling strategies (random and driver-level separation) were employed to assess whether the observed RF classifier performance generalizes across subjects and settings. We reported the average performance $\mathbf{a,P,R,}$ and $\mathbf{F1}$ under two complementary sampling strategies:
Algorithm 1 Scenario-Based Video Analysis
1:Input: Video samples $\mathcal{V}=\{V_{k}^{(D_{i},Sc_{j})}\}$, subjects $\mathcal{D}$, scenarios $\mathcal{S}$, vision model $F_{\theta}$, PCA $g:\mathbb{R}^{m}\!\to\!\mathbb{R}^{n}$, sampling method $\texttt{S}\in\{\text{Random},\text{DLS}\}$
2:Output: Scenario ranking $\{\mathrm{Dist}(Sc_{j})\}_{Sc_{j}\in\mathcal{S}}$, trained RFs $\{\mathbf{a},\mathbf{P},\mathbf{R},\mathbf{F1}\}$
3:for each scenario $Sc_{j}\in\mathcal{S}$ do
4:  $\mathcal{E}_{Sc_{j}}\leftarrow\emptyset$
5:  for each subject $D_{i}\in\mathcal{D}$ do
6:   for each video $V_{k}^{(D_{i},Sc_{j})}$ do
7:     $Z_{k}\leftarrow F_{\theta}\!\big{(}V_{k}^{(D_{i},Sc_{j})}\big{)}$ $\triangleright$ Step 1
8:     $\mathcal{E}_{Sc_{j}}\leftarrow\mathcal{E}_{Sc_{j}}\cup\{Z_{k}\}$
9:   end for
10:  end for
11:  $\tilde{\mathcal{E}}_{Sc_{j}}\leftarrow g(\mathcal{E}_{Sc_{j}})$ $\triangleright$ Step 2
@@
12:  Divide $\tilde{\mathcal{E}}_{Sc_{j}}$ into class-wise sets $\tilde{Z}_{\text{normal}}^{(Sc_{j})},\tilde{Z}_{\text{MD}}^{(Sc_{j})}$
13:  $\boldsymbol{\mu}_{\text{normal}}^{(Sc_{j})}\!\leftarrow\!\frac{1}{|\tilde{Z}_{\text{normal}}^{(Sc_{j})}|}\sum_{\mathbf{z}\in\tilde{Z}_{\text{normal}}^{(Sc_{j})}}\mathbf{z}$, $\boldsymbol{\mu}_{\text{MD}}^{(Sc_{j})}\!\leftarrow\!\frac{1}{|\tilde{Z}_{\text{MD}}^{(Sc_{j})}|}\sum_{\mathbf{z}\in\tilde{Z}_{\text{MD}}^{(Sc_{j})}}\mathbf{z}$
14:  $\mathrm{Dist}(Sc_{j})\leftarrow\left\|\boldsymbol{\mu}_{\text{normal}}^{(Sc_{j})}-\boldsymbol{\mu}_{\text{MD}}^{(Sc_{j})}\right\|_{2}$ $\triangleright$ Step 3
15:end for
[…]
18:for each $Sc\in\{Sc_{\max},Sc_{\min}\}$ do
19:  $(\mathbf{X}^{(Sc)},\mathbf{y}^{(Sc)})\leftarrow(\tilde{\mathcal{E}}_{Sc},\text{labels})$
20:  if $\texttt{S}=\text{Random}$ then
21:   Stratified train/val/test split on samples (balanced)
22:  else if $\texttt{S}=\text{DLS}$ then
23:   Leave-$k$-drivers-out: hold out $k$ drivers as test set; repeat $r$ runs with different held-out drivers
24:  end if
25:  Train RF $\tilde{H}^{(Sc)}:\mathbb{R}^{n}\to\mathcal{Y}$ on train split $\triangleright$ Step 4
26:  Evaluate on held-out data to obtain $\{\mathbf{a},\mathbf{P},\mathbf{R},\mathbf{F1}\}_{Sc}$; average over runs if DLS
27:end for
28:return $\{\mathrm{Dist}(Sc_{j})\}_{Sc_{j}\in\mathcal{S}},\{\mathbf{a},\mathbf{P},\mathbf{R},\mathbf{F1}\}_{Sc_{\max}},\{\mathbf{a},\mathbf{P},\mathbf{R},\mathbf{F1}\}_{Sc_{\min}}$
@@
    Driver-level Separation (DLS): To evaluate the generalizability of the proposed framework across different individuals, we employed a leave-$k$-drivers-out cross-validation strategy [40], where $k$ drivers were excluded from the training set and reserved exclusively for testing. This procedure was repeated over $r$ independent runs, each with a different subset of left-out drivers. Finally, the results across these runs were averaged to get the final performance metrics $\mathbf{(a,P,R,F1)}$. It ensures robustness and mitigates variance due to subject-specific biases.
This sampling strategy evaluates both intra-subject consistency (random) and inter-subject generalization (DLS) in the proposed modeling approach.
@@ 3.1 Scenario-based Roadway Selection
To investigate the connection between driving behavior and cognitive status, we tested and selected two distinct scenarios from the Omaha most transited route where older drivers exhibited the most discriminative driving patterns. These patterns can serve as interpretable behavioral cues for identifying cognitive decline. The selected driving scenarios are:
@@
From different federal functional categories (e.g., Local, Major collector, Minor collector, etc.), we selected the most contrastive scenario combination and performed scenario-specific video sampling as section 2.3. Figure 6 shows the GIS map highlighting the two driving scenarios for our analysis. We considered a single drive having $\leq 0.31$ miles or $\leq 500$ meters in length as a route segment. Omaha most transited route database consists of 24,106 route segments. In total, we filtered out $964$ unique route segments, consisting of $632$ freeway-interchange segments (highlighted in green on Figure 6(b)) and $332$ interstate segments (highlighted in blue on Figure 6(a)). Table 2 summarizes the key statistics – a balance of the selected data and driving exposure of the cognitive groups. Additionally, we selected the route segments that cover at least a minimum number of drives for both cognitive groups. To interpret the driving exposure data distribution in Table 2, we defined several key variables below:
@@
    $C_{u}$: Number of unique trips by the subjects with MCI
@@
    $M_{u}$: Number of unique trips by the subjects with MD
@@
    $T_{u}=N_{u}+C_{u}+M_{u}$
    It does not include repeated trips which means multiple driving sessions recorded on the same route segment by the same individuals
@@
    $T_{rep}$: The average number of times each unique trip was repeated per subject on a route segment, defined as $T_{rep}=\frac{\overline{T}_{u}}{T_{u}}$. It gives an approximation of how often participants repeated drives on a certain route segment. For example, each unique interstate trip was repeated 4.41 times, compared to 3.55 times for freeway interchange trips. Therefore, participants tended to repeat trips on interstate route segments more frequently than on the freeway interchanges.
Table 2: Driving exposure in the route segments for two scenarios
Scenario Normal MCI MD Unique Recurring Recurring trip
categories trips($N_{u}$) trips($C_{u}$) trips($M_{u}$) trips($T_{u}$) trips($\overline{T}_{u}$) factor($T_{rep}$)
@@
$N_{u}$: unique trips by Normal group; $C_{u}$: unique trips by MCI group; $M_{u}$: unique trips by MD group; $T_{u}$: total unique trips; $\overline{T}_{u}$: total non-unique trips; Fwy-int.: Freeway interchange.
We leveraged “unique drives” vs. “repeated drives” per subject across all the route segments to indicate statistical reliability and diversity in driving exposure by the two cognitive groups. The recurring trips per subject suggest that each participant had substantial driving trips with the selected roadway scenario, offering a significant amount of video data for our analysis.
Figure 7 shows a boxplot distribution of speed limits and the trip distance for both scenarios. As we observed, the freeway-interchange segments tend to have lower speed limits and shorter distances, which work as a transitional driving scenario, involving challenging driving maneuvers and fast decision-making. In contrast, the interstate segments exhibit higher speed limits and longer travel distances, which indicate an uninterrupted highway driving scenario with relaxed driving behavior.
Figure 7: Boxplot distributions of the speed limit and trip distance for Interstate and Freeway-interchange roadways. The middle line shows the median speed limit and median route distances.
@@
    Fwy-interchange scenario, $S_{\text{fwy}}=37.5$ mph
@@
    Fwy-interchange scenario, $D_{\text{fwy}}=0.12$ miles